Journalist Report – February 2nd

Crew204 Journalist Report 02-FEB-2019

Title: Robotic Manipulation
Author: Sonal Baberwal

In today’s world, robots seem to play an important role in our day to day life but the true fact is, for future Space exploration, the robotic technology will play a crucial role for enhancing the overall performance of the respective mission. The combined efforts of human-robot interaction is the new future! Robots are not living things but perform better than living beings in terms of precision and accuracy. We have already noticed the marvels of Space Robotics in the International Space Station or Planetary exploration in our solar system like rovers on Moon and Mars.

It has been a busy day where everyone had been involved in their individual research and responsibilities. From morning the crew had been working in different labs to collaborate their work like the Science, Engineering and the GreenHab.

Kunal Naik (Space Bot) had been busy disassembling and assembling the parts of the child rover to get it ready and operational for the EVA. The rover seems to be ready to crawl on the Martian terrains.

Commander was able to set up the Arduino system for the 3D printer in the morning. He has been experimenting with different soil samples while preparing different compositions from red soil, Hab garden soil and some Martian soil along with charcoal. His objective is to investigate the capacity of various soil combination that can possibly bring life in the new soil. He had also been busy working on setting up 3D printer’s mechanical assembly for his experiments.

Cosmos Patil had been working for continuous scheduled watering on grenhab plants and the sprouts he had planted.

This was followed by the EVA. the participants of the EVA today are Cosmos Patil and Kunal Naik. The had explored the white moon. They had an amazing experience while climbing white moon. A quick test of the child rover had been done and the major purpose was to take photos from child rover but had a system failure.

Kunal Naik has been working over the troubleshooting with the help of Sipna College of Engineering and Technology, India. Tomorrow will be a new day with the hope that our rover crawls one again!

Operations Report – February 2nd

Crew 204 Operations Report 02- FEB – 2019

SOL: 5

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1630 (2nd Feb), Stopped by this morning 0830 (2nd Feb)

Solar— SOC % – Turned on (1st Feb) 72%, Turned off (2nd Feb) 72%; Turned on (2nd Feb) 55%

Diesel Reading – 42%

Propane Reading – 69%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 75%, ~ 375 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 243 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01407339 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Spirit rover used: Used for EVA

Hours: 72.1 hours

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Opportunity rover used: Used for the EVA

Hours: 47.6 hours

Beginning charge: 100%

Ending charge: 65%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: Not Used

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Not Used

Notes on rovers: None

ATV’s Used: None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? Used to go to the Town

Crew Car used and why, where? Off Campus

General notes and comments: Need to take 2 rovers as the EVA was long.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Used some engineering tools for Rover and 3D clay printer project.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: May need a quick removal tissue box.

Greenhab Report – February 2nd

Crew 204 Green Hab Report – 2-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.8°C

High: 25.6°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 243 gallons

Time(s) of watering for crops:


1) 08.45 AM 18.5° C / 41 %
2) 12.01 AM 19.3° C / 33 %
3) 15:00 PM 20.2° C / 38 %
4) 18:00 PM 19.3° C / 41 %

Changes to crops: None

Narrative: Today we observed planted sprouts, their germination is happening as expected in the soil. We made various compositions of soils from collected soils in MDRS, JSC Mars one (artificial soil), garden soil and charcoal. We are going to do an experiment on various compositions prepared by these soils.

Harvest: 1) Cilantro 5 gram
2) Green Chord 20 gram
3) Mixed Salads 5 gram
4) Baby Greens 6 gram

Support/supplies needed: None

Sol Summary – February 2nd

Crew 204 Sol Summary Report 02- FEB – 2019

Sol: 5

Summary Title: The Robotic Manipulation

Author’s name: Kunal Naik (Space_bot)

Mission Status: Nominal

Sol Activity Summary: Robots are the first thing that are sent to the unexplored territories of the planet to reveal the unknown. They are not the living things but does the work better than the living beings. The crew woke up as per the schedule assigned earlier, concluding the morning activities and breakfast with a round table discussion over today’s individual task. The crew engineer (Space Bot) started to work on his child rover’s project. The rover was dis-integrated for system check at the RAMM operations and was again assembled, integrated and tested by generating a local server. The quick test was done during the EVA. While Crew Commander (Carbon Flight) was busy working on his 3D clay printer. The printer was assembled, and Arduino was programmed but integration is yet to be done. The project was assisted by Crew Health and Safety Officer (Mars Bag). The GreenHab Officer (Cosmos) was into his daily assigned Green Hab activities. Lunch (Red Kidney beans with rice) was prepared by Cosmos after his routine Green Hab activites. Later, the crew initiated the PreEVA procedures. The EVA was carried by Cosmos and SpaceBot, the purpose of EVA was to test the child rover and Martian Photography via child rover (which was not successful). Post EVA the team started preparing for Dinner prior to the Sunset, opening the floor for personal opinions, suggestions and recreational activites.

Look Ahead Plan: The 3D clay printer will be integrated by next day, proceeding with testing and printing manually out of different soil samples. The work on the child rover will be to check the Audio/Video system plus checking the battery life for further operations. The child-rover will be operated for future EVA’s. The sprouts and mix seeds are already planted in the GreenHab, hoping to get good results.

Anomalies in work: None Encountered

Weather: Partially Cloudy

Crew Physical Status: Nominal

EVA’s location: 517250 E, 4254300 N (The White Moon region)

Reports to be file: None

Support Requested: May require quick removal tissues.

Copyright © The Mars Society. All rights reserved. | Main Site